OpenStudy (kkrpbra):

In February 2010, about 39.7 million Americans were receiving food stamps. This number is a 22% increase from the same month the year before. Based on these numbers, how many Americans were receiving food stamps in February 2009? Choose the closest answer.

OpenStudy (inkyvoyd):

39.7*(1.00-0.22) is what you are looking for.

OpenStudy (anonymous):

= 39.7 / 1.22 = 32.55 mil

OpenStudy (anonymous):

\[39.7 = x_{2009}*1.22\] \[x_{2009} = 39.7/1.22\] \[x_{2009} = 32.541\] How many sig fig/d.p.?